A 5.38 kg object falls freely (ignore air resistance), after being dropped from rest. Determine the final kinetic energy (in J) for the first meter of the fall.
1
Expert's answer
2019-04-02T10:45:53-0400

Let's first find the velocity of the object at the end of the first meter of the fall from the kinematic equation:


"v_f^2 = v_0^2 + 2gs,"

here, "v_0 = 0" is the initial velocity of the object (at rest), "v_f" is the velocity of the object at the end of the first meter of the fall, "g" is the acceleration due to gravity and "s" is the distance traveled by the object.

Then, we get:


"v_f = \sqrt{2gs} = \sqrt{2 \cdot 9.8 \dfrac{m}{s^2} \cdot 1 m} = 4.43 \dfrac{m}{s}."

Finally, from the definition of the kinetic energy, we can calculate the final kinetic energy for the first meter of the fall:


"KE_f = \dfrac{1}{2}mv_f^2 = \dfrac{1}{2} \cdot 5.38 kg \cdot (4.43 \dfrac{m}{s})^2 = 52.8 J."

Answer:

"KE_f = 52.8 J."